OpenStudy (anonymous):

Find the summation from i = 0 to n of 2^i. Any links that are provided is appreciated

OpenStudy (experimentx):

what's the formula of geometric sequence sum?

OpenStudy (anonymous):

(1-r^n)/1-r. This is not a fraction, is it still geometric series?

OpenStudy (experimentx):

use it ... the this is geometric series. with r=2, a = 1

OpenStudy (experimentx):

just change this into (r^n-1)/r-1

OpenStudy (anonymous):

So then the answer is 1-2^n / 1 - 2 = -(1-2^n)?

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=Sum [2^i%2C+{i%2C+0%2C+10}]+%3D+2^11+-+1

OpenStudy (experimentx):

including i=zero, you have n+1 terms.

OpenStudy (anonymous):

ok

OpenStudy (experimentx):

http://en.wikipedia.org/wiki/Geometric_series#Formula

OpenStudy (anonymous):

Thanks, I'm aware of the formula, I just don' t see how wolfram alpha gets their answer

OpenStudy (experimentx):

I get's it's anwer by adding!! it's just a computer.

OpenStudy (anonymous):

There is no final answer since the answer contains n

OpenStudy (experimentx):

W|A doesn't give you answer in that case. your answer is simply \[ 2^{n+1} - 1 \]

OpenStudy (anonymous):

Right, and if I plug that into the gemoetric series that is not what I get

OpenStudy (experimentx):

Yeah ..

OpenStudy (anonymous):

So the answer varies with what wolfram alpha yields

OpenStudy (experimentx):

No ... just WA does numerical calculation for you.

OpenStudy (anonymous):

I don't care about the numerical calculation

OpenStudy (experimentx):

W|A gives you the same result http://www.wolframalpha.com/input/?i=Sum%5B2%5Ei%2C+%7Bi%2C+0%2C+n%7D%5D

OpenStudy (anonymous):

When I to it myself I get 1-(2^n)/-1 That is for when I plug it into the formula. WA differs obviously

OpenStudy (experimentx):

change that n into n+1

OpenStudy (anonymous):

Why?

OpenStudy (experimentx):

see the formula ... that formula is for n terms ... you have n+1 terms.

OpenStudy (anonymous):

The summation is from i = 0 to n. I do not see a n+1

OpenStudy (anonymous):

Unless you said that backwards

OpenStudy (experimentx):

1,2, ....,n <-- these are n terms include 0 <-- this makes n+1 terms.

OpenStudy (anonymous):

WA has a positive result, mine is negative

OpenStudy (experimentx):

1-(2^n)/-1 = -1(-(2^n - 1)) ... just multiply it by -1 on both numerator and denominator.

OpenStudy (anonymous):

What let's you multiply by a -1 on the numerator and denominator? Don't you need this to be equal to something in order to do that?

OpenStudy (experimentx):

no .. you can always do that.

OpenStudy (anonymous):

hm. ok, thank you

OpenStudy (experimentx):

yw ... and change that n to n+1 if you are summing from i=0 to i=n

OpenStudy (anonymous):

Got it